# PharmacoMatch: Efficient 3D Pharmacophore Screening through Neural Subgraph Matching

### v1

@@ 4 Methodology
The PharmacoMatch model correlates the matching of a query and a target pharmacophore with the partial ordering of their vector representations. Positive pairs represent successful matchings, while negative pairs serve as counter examples. In order to create these pairs from unlabeled training data, we define three families of augmentations $\mathcal{T}$, which are composed of random point deletions and positional point displacements.
For positive pairs, valid queries are created by randomly deleting some nodes from a pharmacophore $P$, leaving at least three, and displacing the remaining nodes within a tolerance sphere of radius $r_{T}$. This augmentation, denoted as $t_{1}(\cdot)\sim\mathcal{T}_{1}$, produces the positive pair $(t_{1}(P),P)$.
Negative pairs are used to show the model examples of unsuccessful matching, employing three strategies that illustrate different types of undesired outcomes. Our first strategy provides the model with examples of positional mismatch, by placing the pharmacophoric points of $P$ on the surface of the tolerance sphere without any point deletions. This augmentation, denoted by $t_{2}(\cdot)\sim\mathcal{T}_{2}$, is used to generate the negative query-target pair $(t_{2}(P),P)$.
Our second strategy teaches the model that every pharmacophoric point in the query should correspond to a point in the target. This is achieved by deleting some target nodes, using an augmentation operator $t_{3}(\cdot)\sim\mathcal{T}_{3}$, where $\mathcal{T}_{3}$ involves node deletion without displacement. As a result, the query in the pair $(t_{1}(P),t_{3}(P))$ only partially matches its target.
With the third strategy, we train the model to avoid matching queries with targets that are significantly different. This approach involves randomly mapping queries $t_{1}(P_{i})$ to the incorrect targets $P_{j}$, where $i\neq j$ (for more details, see Appendix A.2).
Curriculum learning
We design a curriculum learning strategy for training on pharmacophore graphs. We start training with pharmacophores containing four nodes. If the loss does not decrease significantly within 10 epochs, we add pharmacophores with one additional node to the training data. This approach allows the model to start with very simple examples, gradually increasing the difficulty of the matching task.
Model Training
@@
Our GNN encoder model is implemented with three convolutional layers with an output dimension of 64. The MLP has a depth of three dense layers with a hidden dimension of 1024 and an output dimension of 512. The final model was trained for $500$ epochs using an Adam optimizer with a learning rate of $10^{-3}$. The margin of the best performing model was set to $\alpha=100$. The tolerance radius $r_{T}$ for node displacement was set to $1.5$, which is the default value in the pharmacophore screening functionalities of the CDPKit (see Appendix A.5 for more details).
@@
We are using the trained GNN encoder $f_{\Theta}$ to precompute vector embeddings $\mathbf{z}_{T}$ of the database pharmacophores. These are queried with the pharmacophore embedding $\mathbf{z}_{Q}$ by verification of the partial ordering constraint (3), which shall not be violated by more than a threshold $t$. This leads to the decision function $g:\mathbb{R}^{D}_{+}\times\mathbb{R}^{D}_{+}\rightarrow\{0,1\}$:
@@
which evaluates to $1$ if the partial ordering on $\mathbf{z}_{Q}$ and $\mathbf{z}_{T}$ reflects a pharmacophore matching, and 00 otherwise. In the following, we will refer to equation (4) as matching decision function. In practice, we recommend a decision threshold of $t=6500$, which was determined during our benchmark experiments.
@@ 5 Experiments
We designed our embeddings to reflect the type and relative positioning of pharmacophoric points. Comparison of embedding vectors via the matching prediction function should emulate the matching of the underlying pharmacophores. To get a better understanding of the encoder’s latent space, we investigate these properties as follows:
@@
    Pharmacophoric point perception: We investigate the learned embedding space quantitatively through dimensionality reduction.
@@
    Virtual screening performance: The performance of our model is evaluated using ten DUD-E targets, and the produced hitlists are compared with the performance and runtime of the CDPKit (Seidel 2024) alignment algorithm.
DUD-E benchmark dataset
We perform our experiments on the DUD-E benchmark dataset (Mysinger et al. 2012), which is commonly used to evaluate the performance of molecular docking and structure-based screening. The complete benchmark contains 102 protein targets, each accompanied by active and decoy ligands in the form of SMILES strings (Weininger 1988) and the PDB template (Burley et al. 2017) of the ligand-receptor complex. We randomly select ten different protein targets for the evaluation of our model. Ligands in these datasets are processed according to the data curation pipeline outlined in the Methodology section, except that we sample up to 25 conformations per compound. The ligand-receptor complex is used to generate a structure-based query with 5-7 pharmacophoric points (see Appendix A.6 for more details).
5.1 Pharmacophoric point perception
We conduct a quantitative analysis through dimensionality reduction to gain a first intuition for the properties of the learned embedding space.
The partial ordering of graph representations in the embedding space, based on the number of nodes per graph, is essential for encoding query-target relationships. This ordering property of the embedding space can be visualized using principal component analysis (PCA). Figure 5a displays the first two principal component axes of the learned representations, with the representations labeled according to the number of pharmacophoric points of the corresponding pharmacophore. This visualization demonstrates how the embedding vectors are systematically ordered relative to the number of nodes in each pharmacophore graph.
Similarly, the Uniform Manifold Approximation and Projection (UMAP) algorithm (McInnes, Healy, and Melville 2020), a dimensionality reduction technique that preserves the local neighborhood structure of high-dimensional data, was employed. Figure 5b shows the UMAP representation of the embeddings, labeled by the number of pharmacophoric points of a specific type. This visualization suggests that pharmacophores with a similar set of points are mapped proximally within the order embedding space.
Figure 5: (a.) Dimensionality reduction of the ADA target’s embedding space via PCA, with embeddings labeled by pharmacophoric point count. (b.) Dimensionality reduction via UMAP, with embeddings labeled by pharmacophoric point type. (c.) Experimental validation of the model’s perception of 3D point positions, showing the mean matching decision function versus the displacement radius $r_{D}$ of the augmentation, with a decision threshold set to $t=6500$.
@@ 5.2 Positional perception
We define a family of augmentations $\mathcal{T}_{r_{D}}$ to randomly delete nodes from a pharmacophore $P$ and displace the remaining nodes by a radius $r_{D}$. We sample augmentations $t_{r_{D}}(\cdot)\sim\mathcal{T}_{r_{D}}$ with increasing radius $r_{D}$ taken from a uniform grid of $m$ distances between 0 and 10 Å.
For a given batch of pharmacophores $\{P_{1},...,P_{n}\}$, we generate the query-target pairs $\{(t_{r_{D}}(P_{1}),P_{1}),...,(t_{r_{D}}(P_{n}),P_{n})\}$. We then evaluate the decision function $g(\cdot,\cdot)$ (Equation 4) on the corresponding vector representations and calculate the mean of the decision function across all pairs against an increasing radius $r_{D}$, which is illustrated in Figure 5c.
Without node displacement, the mean matching decision function is close to 1, indicating that the model recognizes pharmacophores with reduced node sets as valid queries. With a displacement of approximately 1.5 Å, the mean matching decision value drops to 50%, demonstrating the model’s consideration of the chosen tolerance radius. Beyond a displacement of 1.5 Å, the decision function further decreases, approaching a plateau at approximately 6 Å. The results show that our model integrates 3D-positional information of pharmacophoric points into the learned representations.
5.3 Virtual screening
@@
Baseline algorithm
The baseline for our comparison is the alignment algorithm implemented in the open-source software CDPKit (Seidel 2024), which utilizes clique-detection followed by Kabsch alignment (Kabsch 1976). The alignment of a query $P_{Q}$ and a target $P_{T}$ is evaluated with an alignment score $S:\mathcal{P}\times\mathcal{P}\rightarrow\mathbb{R}_{+}$, which takes into account the number of matched features and their geometric fit (further details are provided in the Appendix A.6). The ligand ranking score is calculated as $\psi_{i}=\max(\{S(P_{Q},P_{1}),...,S(P_{Q},P_{k_{i}})\}_{i})$, the highest alignment score represents the score for the database ligand. Analogous to equation (4), we can also define a matching decision function $\phi$ based on the alignment score, where $t=|P_{Q}|$:
@@
Evaluation
Both algorithms rank database ligands to produce a hitlist. We assess the performance of PharmacoMatch on the benchmark using two approaches. First, we demonstrate that the PharmacoMatch penalty $E(\cdot,\cdot)$ correlates with the matching decision function $\phi(\cdot,\cdot)$ of the alignment algorithm. We evaluate both functions against all pharmacophores in a dataset w.r.t. query $P_{Q}$. The outputs are compared by generating the corresponding receiver operating characteristic (ROC) curves, and the performance is quantified using the area under the ROC curve (AUROC) metric.
@@
Second, we compare the virtual screening performance of our model and the alignment algorithm using the ligand ranking score $\psi$. The primary objective of virtual screening is to find active compounds amongst decoys. We evaluate this using two different metrics. The AUROC metric is used to evaluate the overall classification performance w.r.t. activity label $y_{i}$. A drawback of this metric is that it does not reflect the early enrichment of active compounds in the hitlist, which is of significant interest in virtual screening. Early enrichment is assessed using the Boltzmann-enhanced discrimination of ROC (BEDROC) metric (Truchon and Bayly 2007), which assigns higher weights to better-ranked samples. Note that these performance metrics are entirely dependent on the chosen query. Rather than aiming to maximize those metrics, our goal is to achieve comparable values between our model and the alignment algorithm.
Screening performance
Our results, comparing PharmacoMatch with the alignment algorithm across all ten targets, are summarized in Table 1 (ROC plots are provided in the Appendix A.6). We observe a robust correlation between the hitlists generated by the two algorithms, demonstrating the effectiveness of our approach. This correlation varies by target, reflecting the sensitivity of virtual screening to the chosen query. Although the alignment algorithm achieves generally higher AUROC scores and early enrichment, our method consistently produces hitlists with competitive performance across several targets.
@@
Table 1: Method comparison and screening performance of the PharmacoMatch algorithm and the CDPKit alignment algorithm on ten different DUD-E protein targets (see Appendix A.6 for details). BEDROC values are calculated with $\alpha=20$, as recommended by Truchon and Bayly (2007). Metrics are reported in percent. Confidence intervals are calculated using bootstrapping (Efron 1979), with standard deviations reported based on 1,000 resampled datasets.
Protein Method Screening Performance
Target Comparison PharmacoMatch CDPKit
AUROC AUROC BEDROC AUROC BEDROC
ACES $90.7\pm 0.2$ $57\pm 2$ $19\pm 2$ $55\pm 1$ $15\pm 1$
ADA $97.7\pm 0.3$ $80\pm 3$ $39\pm 4$ $93\pm 2$ $80\pm 4$
ANDR $98.0\pm 0.2$ $78\pm 2$ $32\pm 2$ $72\pm 2$ $26\pm 2$
EGFR $90.4\pm 0.5$ $59\pm 1$ $10\pm 1$ $72\pm 1$ $22\pm 1$
FA10 $84.2\pm 0.1$ $48\pm 1$ $2.1\pm 0.4$ $55\pm 1$ $5.8\pm 0.6$
KIT $80.2\pm 0.1$ $52\pm 2$ $2.3\pm 0.6$ $58\pm 2$ $7\pm 1$
PLK1 $79.1\pm 0.5$ $61\pm 3$ $9\pm 2$ $75\pm 3$ $39\pm 3$
SRC $95.6\pm 0.1$ $77\pm 1$ $22\pm 1$ $77\pm 1$ $24\pm 1$
THRB $86.4\pm 0.4$ $73\pm 1$ $26\pm 2$ $81\pm 1$ $40\pm 2$
UROK $83.3\pm 0.2$ $59\pm 2$ $3\pm 1$ $91\pm 1$ $52\pm 3$
Runtime comparison
@@
In terms of runtime, PharmacoMatch significantly outperforms the alignment algorithm. We compare the time required for alignment, embedding, and vector matching per pharmacophore. Alignment is performed in parallel on an AMD EPYC 7713 64-Core Processor with 128 threads, while pharmacophore embedding and matching are run on an NVIDIA GeForce RTX 3090, with both devices having comparable purchase prices and release dates. Creating vector embeddings from pharmacophore graphs takes $92\pm 12~{}\mathrm{\mu s}$ per pharmacophore, which takes longer than aligning a query to a target with $13\pm 7~{}\mathrm{\mu s}$. However, the embedding process only needs to be performed once. Subsequently, the preprocessed vector data can be used for vector matching, which takes $0.30\pm 0.09~{}\mathrm{\mu s}$, being approximately two orders of magnitude faster than the alignment. Additionally, vector comparison is independent of the query size, an advantage not shared by the alignment algorithm. Although executed on different hardware, this comparison highlights the speed-gain of our algorithm.
Practical considerations
@@
There are two options for integrating our model into a virtual screening pipeline. First, the PharmacoMatch model can be used in place of the alignment algorithm to generate a hitlist of ligands, which is suitable for quickly producing a compound list for experimental testing. Alternatively, our method can serve as an efficient prefiltering tool for very large databases, reducing the number of molecules from billions to millions, after which the slower alignment algorithm can be applied to this filtered subset. Note that alignment will still be necessary if visual inspection of aligned pharmacophores and corresponding ligands is desired.

### v2

@@ 4 Methodology
The PharmacoMatch model correlates the matching of a query and a target pharmacophore with the partial ordering of their vector representations. Positive pairs represent successful matchings, while negative pairs serve as counter examples. In order to create these pairs from unlabeled training data, we define three families of augmentations $\mathcal{T}$, which are composed of random point deletions and positional point displacements. For positive pairs, valid queries are created by randomly deleting some nodes from a pharmacophore $P$, leaving at least three, and displacing the remaining nodes within a tolerance sphere of radius $r_{T}$. This augmentation, denoted as $t_{1}(\cdot)\sim\mathcal{T}_{1}$, produces the positive pair $(t_{1}(P),P)$. Negative pairs highlight examples of unsuccessful matching, using three strategies to capture different undesired outcomes. The first strategy introduces positional mismatches by displacing the pharmacophoric feature points of $P$ to the boundary of the tolerance sphere without deleting any points. Specifically, we compute the mean position of the points, $\mathbf{\mu}=\frac{1}{|P|}\sum_{i=1}^{|P|}\mathbf{r}_{i}$, and shift each point $p_{i}$ by $r_{T}$ along the direction $\mathbf{r}_{i}-\mathbf{\mu}$. This approach avoids the unintended creation of positive pairs, which can occur with random sampling, and ensures that displacements do not cancel out. The resulting augmentation, denoted $t_{2}(\cdot)\sim\mathcal{T}_{2}$, generates the negative query-target pair $(t_{2}(P),P)$. This strategy demonstrated better model performance than random positional displacement. Our second strategy teaches the model that every pharmacophoric feature point in the query should correspond to a point in the target. This is achieved by deleting some target nodes, using an augmentation operator $t_{3}(\cdot)\sim\mathcal{T}_{3}$, where $\mathcal{T}_{3}$ involves node deletion without displacement. As a result, the query in the pair $(t_{1}(P),t_{3}(P))$ only partially matches its target. With the third strategy, we train the model to avoid matching queries with targets that are significantly different. This approach involves randomly mapping queries $t_{1}(P_{i})$ to the incorrect targets $P_{j}$, where $i\neq j$ (for more details, see Appendix A.2).
Model details & design choices
@@
Our GNN encoder model is implemented with three convolutional layers with an output dimension of 64. The MLP has a depth of three dense layers with a hidden dimension of 1024 and an output dimension of 512. The final model was trained for $500$ epochs using an Adam (Kingma, 2014) optimizer with a learning rate of $10^{-3}$. The margin of the best performing model was set to $\alpha=100$. The default tolerance radius $r_{T}$ in CDPKit’s pharmacophore screening is set to 1.5 Å, and we use the same value for the node displacement during model training to ensure consistency with the alignment algorithm in subsequent evaluations. We design a curriculum learning strategy for learning on pharmacophore graphs, detailed in Appendix A.5, along with details on model training and hyperparameter optimization. Systematic ablation studies of model architecture components and the impact of different augmentation strategies, model size, and embedding dimension on model performance are provided in Appendix A.5.
@@
We use the trained GNN encoder $f_{\Theta}$ to precompute vector embeddings $\mathbf{z}_{T}$ of the database pharmacophores. These are queried with the pharmacophore embedding $\mathbf{z}_{Q}$ by verification of the partial ordering constraint through penalty function $E(\cdot,\cdot)$ (Equation 1), which shall not exceed a threshold $t$. This leads to decision function $g:\mathbb{R}^{D}_{+}\times\mathbb{R}^{D}_{+}\rightarrow\{0,1\}$:
@@
evaluating $1$, if the partial ordering on $\mathbf{z}_{Q}$ and $\mathbf{z}_{T}$ reflects a pharmacophore matching, and 00 otherwise. In the following, we refer to equation (2) as matching decision function. In practice, we recommend a decision threshold of $t=6500$, determined from our benchmark experiments.
@@ 5 Experiments
We designed the embeddings to reflect the type and relative positioning of pharmacophoric feature points. Comparison of embedding vectors via the matching prediction function should emulate the matching of the underlying pharmacophores. To get a better understanding of the encoder’s latent space, we investigate these properties as follows:
@@
    Pharmacophoric feature point perception: We investigate the learned embedding space qualitatively through dimensionality reduction.
@@
    Virtual screening performance: The performance of our model is evaluated using ten DUD-E (Mysinger et al., 2012) targets, and the produced hitlists are compared with the performance and runtime of the CDPKit (Seidel, 2024) alignment algorithm. We further evaluate the pre-screening performance of our model and compare it against the PharmacoNet (Seo & Kim, 2023) model using the DEKOIS2.0 (Bauer et al., 2013) and LIT-PCBA (Tran-Nguyen et al., 2020) benchmark datasets.
Benchmark datasets
We perform experiments on the DUD-E benchmark dataset (Mysinger et al., 2012), which is commonly used to evaluate the performance of molecular docking and structure-based screening. The complete benchmark contains 102 protein targets, each accompanied by active and decoy ligands in the form of SMILES strings (Weininger, 1988) and the PDB template (Burley et al., 2017) of the ligand-receptor complex. We randomly select ten different protein targets for a proof-of-concept comparison with the CDPKit alignment algorithm. For our pre-screening experiment, we use the DEKOIS2.0 (Bauer et al., 2013) dataset, which contains 80 targets, each with 40 actives and 1,200 decoys, as well as the LIT-PCBA (Tran-Nguyen et al., 2020) dataset, consisting of 15 target sets with 7,761 confirmed actives and 382,674 inactive compounds. Database ligands are processed according to the data curation pipeline outlined in the Methodology section, with the exception that we sample up to 25 conformations per compound. Further information is provided in the Appendix A.6.
5.1 Pharmacophoric feature point perception
We conduct a qualitative analysis through dimensionality reduction to gain a first intuition for the properties of the learned embedding space. The partial ordering of graph representations in the embedding space, based on the number of nodes per graph, is essential for encoding query-target relationships. This ordering property of the embedding space can be visualized using principal component analysis (PCA). Figure 4a displays the first two principal component axes of the learned representations, with the representations labeled according to the number of pharmacophoric feature points of the corresponding pharmacophores. This visualization demonstrates how the embedding vectors are systematically ordered relative to the number of nodes in each pharmacophore graph. Similarly, the Uniform Manifold Approximation and Projection (UMAP) algorithm (McInnes et al., 2020), a dimensionality reduction technique that preserves the local neighborhood structure of high-dimensional data, was employed. Figure 4b shows the UMAP representation of the embeddings, labeled by the number of feature points of a specific type. This visualization suggests that pharmacophores with a similar set of points are mapped proximally within the embedding space.
@@ 5.2 Positional perception
We define a family of augmentations $\mathcal{T}_{r_{D}}$ to randomly delete nodes from a pharmacophore $P$ and displace the remaining nodes by a radius $r_{D}$. We sample augmentations $t_{r_{D}}(\cdot)\sim\mathcal{T}_{r_{D}}$ with increasing radius $r_{D}$ taken from a uniform grid of $m$ distances between 0 and 10 Å. For a given batch of pharmacophores $\{P_{1},...,P_{n}\}$, we generate the query-target pairs $\{(t_{r_{D}}(P_{1}),P_{1}),...,(t_{r_{D}}(P_{n}),P_{n})\}$. We then evaluate the decision function $g(\cdot,\cdot)$ (Equation 2) on the corresponding vector representations and calculate the mean of the decision function across all pairs against an increasing radius $r_{D}$, which is illustrated in Figure 4c. Without node displacement, the mean matching decision function is close to 1, indicating that the model recognizes pharmacophores with reduced node sets as valid queries. With a displacement of approximately 1.5 Å, the mean matching decision value drops to 50%, demonstrating the model’s consideration of the chosen tolerance radius. Beyond a displacement of 1.5 Å, the decision function further decreases, approaching a plateau at approximately 6 Å. The results show that our model integrates 3D-positional information of pharmacophoric feature points into the learned representations.
Figure 4: (a.) Dimensionality reduction of the ADA target’s embedding space via PCA, with embeddings labeled by pharmacophoric feature point count. (b.) Dimensionality reduction via UMAP, with embeddings labeled by pharmacophoric feature point type. (c.) Experimental validation of the model’s perception of 3D point positions, showing the mean matching decision function versus the displacement radius $r_{D}$ of the augmentation, with a decision threshold set to $t=6500$.
5.3 Alignment prediction and performance as a pre-screening tool
Hitlist ranking
@@ 5.3 Alignment prediction and performance as a pre-screening tool
Ground truth for alignment prediction
We evaluate the alignment prediction performance of PharmacoMatch by relative comparison with the alignment algorithm implemented in the open-source software CDPKit (Seidel, 2024), which utilizes clique-detection followed by Kabsch alignment (Kabsch, 1976). The ligand-receptor complex of the respective DUD-E target is used to generate a structure-based pharmacophore query with the CDPKit. The CDPKit alignment algorithm only returns exact matches, meaning that queries with an excessive number of points may yield no results. Consequently, pharmacophore modeling often requires user interaction to reduce the number of points in the query. For our comparison, we refined the initial query pharmacophores to a subset of 5–7 pharmacophoric feature points, which is a common range in pharmacophore modeling. These points were selected to ensure that the query yields meaningful enrichment for the CDPKit algorithm. Only after this refinement did we proceed to compare PharmacoMatch against this ground truth. The alignment of a query $P_{Q}$ and a target $P_{T}$ is evaluated with an alignment score $S:\mathcal{P}\times\mathcal{P}\rightarrow\mathbb{R}_{+}$, which takes into account the number of matched features and their geometric fit (further details are provided in the Appendix A.6). The ligand ranking score is calculated as $\psi_{i}=\max(\{S(P_{Q},P_{1}),...,S(P_{Q},P_{k_{i}})\}_{i})$, the highest alignment score represents the score for the database ligand. Analogous to equation (2), we can also define a matching decision function $\phi$ based on the alignment score, where $t=|P_{Q}|$:
@@
Evaluation metrics
@@
Both algorithms rank database ligands to produce a hitlist. We assess the performance of PharmacoMatch on the benchmark using two approaches. First, we demonstrate that the PharmacoMatch penalty $E(\cdot,\cdot)$ correlates with the matching decision function $\phi(\cdot,\cdot)$ of the alignment algorithm. We evaluate both functions against all pharmacophores in a dataset w.r.t. query $P_{Q}$. The outputs are compared by generating the corresponding receiver operating characteristic (ROC) curves, and the performance is quantified using the area under the ROC curve (AUROC) metric. We argue that this relative performance is the key metric for evaluating the alignment prediction performance of our model. Second, we compare the absolute screening performance of our model and the alignment algorithm using the ligand ranking score $\psi$. The primary objective of virtual screening is to find active compounds among decoys. The AUROC metric is used to evaluate the overall classification performance w.r.t. activity label $y_{i}$. A drawback of this metric is that it does not reflect the early enrichment of active compounds in the hitlist, which is of significant interest in virtual screening. Early enrichment is assessed using the enrichment factor (EFα%) and the Boltzmann-enhanced discrimination of ROC (BEDROC) metric (Truchon & Bayly, 2007), which assigns higher weights to better-ranked samples (definitions in Appendix A.6). Note that these performance metrics are entirely dependent on the chosen query. Therefore, our primary performance metric in this experiment is the relative performance AUROC. For completeness, we also report absolute virtual screening metrics. We conduct this proof-of-concept comparison using a randomly selected subset of 10 DUD-E targets, serving as a focused case study.
@@
Table 1: Method comparison and screening performance of the PharmacoMatch algorithm and the CDPKit alignment algorithm on ten different DUD-E protein targets (see Appendix A.6 for details). BEDROC values are calculated with $\alpha=20$, as recommended by Truchon & Bayly (2007), AUROC and BEDROC are reported in percent. Confidence intervals are calculated using bootstrapping (Efron, 1979), with standard deviations reported based on 100 resampled datasets.
Protein Relative Absolute screening performance Absolute screening performance
target performance PharmacoMatch CDPKit
[…]
ACES $96.0\pm 0.2$ $58\pm 2$ $18\pm 1$ $8.4\pm 1.4$ $3.5\pm 0.3$ $2.2\pm 0.2$ $55\pm 1$ $16\pm 2$ $5.5\pm 1.3$ $3.0\pm 0.3$ $2.1\pm 0.2$
ADA $98.3\pm 0.2$ $83\pm 3$ $44\pm 4$ $16.7\pm 4.1$ $9.5\pm 1.0$ $5.7\pm 0.4$ $94\pm 1$ $82\pm 3$ $53.6\pm 4.3$ $15.9\pm 0.9$ $8.4\pm 0.4$
ANDR $98.8\pm 0.1$ $76\pm 1$ $33\pm 2$ $15.8\pm 1.9$ $6.0\pm 0.5$ $4.3\pm 0.3$ $71\pm 2$ $26\pm 2$ $12.6\pm 2.1$ $4.4\pm 0.5$ $3.7\pm 0.3$
EGFR $91.1\pm 0.5$ $63\pm 1$ $11\pm 1$ $3.1\pm 0.7$ $2.0\pm 0.3$ $1.6\pm 0.2$ $76\pm 1$ $26\pm 2$ $12.2\pm 1.6$ $4.6\pm 0.3$ $3.7\pm 0.2$
FA10 $84.3\pm 0.1$ $47\pm 1$ $1\pm 1$ $0.2\pm 0.2$ $0.1\pm 0.1$ $0.2\pm 0.1$ $55\pm 1$ $6\pm 1$ $0.0\pm 0$ 0.0 $0.7\pm 0.2$ $1.2\pm 0.1$
KIT $84.5\pm 0.1$ $56\pm 2$ $4\pm 1$ $0.0\pm 0$ 0.0 $0.4\pm 0.2$ $0.7\pm 0.2$ $63\pm 2$ $9\pm 2$ $1.1\pm 0.8$ $1.2\pm 0.4$ $1.8\pm 0.3$
PLK1 $79.5\pm 0.5$ $62\pm 3$ $9\pm 2$ $1.5\pm 1.3$ $0.7\pm 0.3$ $1.8\pm 0.3$ $75\pm 3$ $39\pm 3$ $5.7\pm 2.3$ $10.2\pm 0.9$ $5.5\pm 0.5$
SRC $96.6\pm 0.2$ $79\pm 1$ $27\pm 1$ $6.0\pm 1.0$ $5.3\pm 0.4$ $4.6\pm 0.2$ $80\pm 1$ $28\pm 1$ $11.1\pm 1.2$ $5.3\pm 0.4$ $4.3\pm 0.2$
THRB $89.9\pm 0.5$ $70\pm 1$ $22\pm 1$ $5.9\pm 1.0$ $4.8\pm 0.4$ $3.3\pm 0.2$ $79\pm 1$ $35\pm 2$ $11.8\pm 1.5$ $7.2\pm 0.4$ $4.5\pm 0.2$
UROK $83.8\pm 0.2$ $60\pm 2$ $4\pm 1$ $0.6\pm 0.7$ $0.5\pm 0.2$ $0.4\pm 0.2$ $91\pm 1$ $55\pm 3$ $24.5\pm 2.8$ $10.4\pm 0.9$ $8.2\pm 0.4$
Alignment prediction performance
@@
Our results, comparing PharmacoMatch with the CDPKit alignment algorithm across the selected targets, are summarized in Table 1 (ROC plots are provided in the Appendix A.6). We observe a robust correlation between the hitlists generated by the two algorithms, demonstrating the effectiveness of our approach. This correlation varies by target, reflecting the sensitivity of virtual screening to the chosen query. Although the alignment algorithm achieves generally higher AUROC scores and early enrichment, our method consistently produces hitlists with competitive performance across several targets. In terms of runtime, PharmacoMatch significantly outperforms the alignment algorithm. We compare the time required for alignment, embedding, and vector matching per pharmacophore. Alignment is performed in parallel on an AMD EPYC 7713 64-Core Processor with 128 threads, while pharmacophore embedding and matching are run on an NVIDIA GeForce RTX 3090, with both devices having comparable purchase prices and release dates. Creating vector embeddings from pharmacophore graphs takes $67\pm 7~{}\mathrm{\mu s}$ per pharmacophore, approximately as long as aligning a query to a target with $66\pm 6~{}\mathrm{\mu s}$. However, the embedding process only needs to be performed once. Subsequently, the preprocessed vector data can be used for vector matching, which takes $0.3\pm 0.1~{}\mathrm{\mu s}$, being approximately two orders of magnitude faster than the alignment. Additionally, vector comparison is independent of the query size, an advantage not shared by the alignment algorithm. Although executed on different hardware, this comparison highlights the speed-gain of our algorithm.
Table 2: Comparison of PharmacoMatch with PharmacoNet (Seo & Kim, 2023) as a pre-screening tool. The reported performance metrics were averaged over all targets of the DEKOIS2.0 and LIT-PCBA datasets. AUROC and BEDROC are given in percent, BEDROC values are calculated with $\alpha=80.5$, as reported in Seo & Kim (2023). PharmacoNets runtime per ligand depends on the number of atoms, here we report their measurement for ligands with 70 heavy atoms. The runtime for PharmacoMatch was calculated as average over all ligands in the benchmark.
DEKOIS2.0 LIT-PCBA Runtime
[…]
PharmacoNet 62.5 12.3 4.4 4.2 2.9 - - - 3.1 - $5.2\cdot 10^{-3}$
PharmacoMatch (ours) 60.9 15.1 5.5 4.9 3.2 57.4 5.0 6.0 3.5 2.2 $\mathbf{3.3\cdot 10^{-6}}$
Applicability as a pre-screening tool
We demonstrated how PharmacoMatch predicts pharmacophore matching and compared its performance with the CDPKit alignment algorithm. A key aspect of this comparison is the role of user interaction in query design, since the alignment algorithm only identifies a match when all pharmacophoric feature points align, and while it is possible to define feature points as optional, this flexibility significantly increases runtime due to the combinatorial explosion of possible query patterns. Consequently, handling larger queries with ten or more points becomes computationally prohibitive. PharmacoMatch operates independently of the query’s size. This characteristic makes PharmacoMatch particularly suitable as a pre-screening tool. By using structure-based queries without refinement step, it can filter extensive datasets effectively, reducing the computational burden before engaging more resource-intensive methods or requiring user interaction. To evaluate this use case, we compare our method with the recent pre-screening tool PharmacoNet (Seo & Kim, 2023), which employs image segmentation to generate pharmacophore queries and a parameterized analytical function for hitlist creation. To replicate an automated pre-screening scenario, we exclude user interaction from our workflow. Specifically, we generate an interaction pharmacophore using CDPKit and use it directly as a query for PharmacoMatch. We assess the pre-screening performance of PharmacoMatch and PharmacoNet on the DEKOIS2.0 (Bauer et al., 2013) and LIT-PCBA (Tran-Nguyen et al., 2020) benchmark datasets. Performance is evaluated using the average AUROC, BEDROC, and EF metrics across all targets (see Appendix A.6 for more details). The results of this comparison are summarized in Table 2. Our evaluation demonstrates that PharmacoMatch outperforms PharmacoNet with respect to early enrichment. Importantly, PharmacoMatch’s runtime is three orders of magnitude faster. Given that the primary goal of pre-screening is rapid and cost-effective filtering before using more computationally expensive methods, we argue that this substantial improvement in runtime makes PharmacoMatch an effective pre-screening tool.
Practical considerations & limitations
@@
There are two options for integrating our model into a virtual screening pipeline. First, the PharmacoMatch model can be used in place of the alignment algorithm to generate a hitlist of potential active compounds, which is suitable for quickly producing a compound list for experimental testing. Second, our method can serve as an efficient pre-screening tool for very large databases, reducing the number of molecules from billions to millions, after which the slower alignment algorithm can be applied to this filtered subset. Note that alignment will still be necessary if visual inspection of aligned pharmacophores and corresponding ligands is desired. An important limitation is that the embedding process is not lossless and might lead to reduced 3D geometric precision, when compared to an alignment algorithm. The employed E(3)-invariant encoder cannot distinguish a pharmacophore from its mirror image, potentially increasing the false positive rate. Addressing these limitations will be part of future work.